OpenStudy (anonymous):

find the wavelength of an electron with a mass of 9.11*10^-31kilogram and a velocity of 5.4*10^6 meters per sec

OpenStudy (anonymous):

check Matter and wave–particle duality (or Louis de Broglie hypothesis)\[\lambda=\frac{ h }{ m·v }\sqrt{1+\frac{ v^2 }{ c^2 }}\]